# Decentralized Parametric Stability Certificates for Grid-Forming Converter Control

### v5

@@ 5.2 Case Study II: Nonlinear circuit model
For the same three-node system as in Section 5.1, we now use Simscape Electrical in MATLAB/Simulink to perform a detailed electromagnetic transients (EMT) simulation based on a nonlinear circuit model of the three-phase network. Additionally, for each GFM VSC, we incorporate the full nonlinear converter models as shown in Fig. 3. Compared to Case Study I, this increases complexity in two key aspects: first, by considering nonlinear models, and second, by accounting for the full network dynamics without the simplifying assumptions introduced in Section 3.1. The simulation results in Fig. 13 demonstrate that a GFM 3 controller satisfying the decentralized stability conditions in 24 and 25 ensures stability even in the presence of nonlinear network and device models.
Figure 10: Stable system response of the block diagram in Fig. 4 with three GFM devices and the simplified network model 19, where the controller of GFM 3 does not satisfy any of the conditions 24, 25 and 26 (no cond.).
Figure 11: Stable system response of the block diagram in Fig. 4 with three GFM devices and the simplified network model 18, where the GFM 3 controller satisfies the conditions in 26 (cond. L1).
Figure 12: Stable system response of the block diagram in Fig. 4 with three GFM devices and the dynamic network model 17, where all GFM controllers are satisfying the conditions in 24 and 25 (cond. DYN). We also display the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and when it is satisfying the conditions in 26 (cond. L1).
Figure 13: Stable system response of the three-node converter system with nonlinear circuit and device models where all GFM controllers are satisfying the conditions in 24 and 25 (cond. DYN). Additionally, we indicate the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and when it is satisfying the conditions in 26 (cond. L1).
In contrast, when the GFM 3 controller satisfies only 26 or does not meet any stability conditions, interactions with the network dynamics can lead to instability. Overall, we conclude that our stability conditions in 24 and 25, which explicitly account for network dynamics, remain effective in a nonlinear circuit scenario, provided the system operates near the nominal point where linearization errors are sufficiently small. However, the stability conditions, designed for a simplified static network, may fail when network dynamics are taken into account.
@@ 6 Conclusion
Future work includes the extension of our dynamic loop-shifting framework to a non-diagonal structure of the matrix blocks $\Gamma_{i}(s)$ with cross-coupling between active and reactive power, thereby eventually reducing conservatism by requesting coupled device controllers. Beyond that, from an application point of view, we envision our stability framework to support the formulation of new grid codes for future power systems.

### v6

@@ 5.2 Case Study II: Nonlinear circuit model
For the same three-bus system as in Section 5.1, we now use Simscape Electrical in MATLAB/Simulink to perform a detailed EMT simulation based on a nonlinear circuit model of the three-phase network. In particular, we now consider more general and realistic line models also including shunt capacitors. Additionally, for each GFM VSC, we incorporate the full nonlinear converter models as shown in Fig. 3 (using average models). Compared to Case Study I, this increases complexity in two key aspects: first, by considering nonlinear models, and second, by accounting for the full network dynamics without the simplifying assumptions in Section 3.1, while also including shunt capacitors. The simulation results for a small-signal load increase in Fig. 15 demonstrate that a GFM3 controller satisfying the decentralized stability conditions in 25 and 26 maintains stability even in the presence of nonlinear network and device models.
Figure 14: Stable system response of the block diagram in Fig. 4 with three GFM devices and the dynamic network model 18, where all GFM controllers are satisfying the conditions in 25 and 26 (cond. DYN). We also display the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and when it is satisfying the conditions in 27 (cond. L1).
In contrast, when the GFM 3 controller satisfies only 27 or does not meet any stability conditions, interactions with the network dynamics can lead to instability. We thus conclude that our stability conditions in 25 and 26, which account for network dynamics, remain effective in a nonlinear circuit scenario including shunt capacitors, provided the system operates near the nominal point where linearization errors are small. However, the stability conditions, designed for a simplified static network, may fail.
Figure 15: Stable system response of the three-bus converter system with nonlinear circuit and device models (including shunt capacitors) where all GFM controllers are satisfying the conditions in 25 and 26 (cond. DYN). Additionally, we indicate the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and the conditions in 27 (cond. L1).
@@ 6 Conclusion
Future work includes the extension of our framework to a non-diagonal structure of the matrix blocks $\Gamma_{i}(s)$, thereby eventually reducing conservatism by requesting coupled device controllers. Beyond that, from an application point of view, we envision our stability framework to support the formulation of new grid codes for future power systems.